# Fairness-Aware Domain Generalization under Covariate and Dependence Shifts

### v1

@@ Appendix D Ablation Studies
9:         $\mathcal{L}_{inv}(\bm{\theta})=\mathcal{L}^{\prime}_{inv}(\bm{\theta})/m$
10:         $\mathcal{L}(\bm{\theta})=\mathcal{L}_{cls}(\bm{\theta})+\lambda_{1}\cdot\mathcal{L}_{inv}(\bm{\theta})$
11:         $\bm{\theta}\leftarrow\bm{\theta}-\eta_{p}\cdot\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$
12:         $\lambda_{1}\leftarrow\text{max}\{[\lambda_{1}+\eta_{d}\cdot(\mathcal{L}_{inv}(\bm{\theta})-\gamma_{1})],0\}$
@@
15:procedure DataAug($\mathbf{x},z,y$)
16:     $\mathbf{c}=E^{c}(E^{m}(\mathbf{x},\bm{\theta}^{m}),\bm{\theta}^{c})$
@@
19:     $\mathbf{x}^{\prime}=G^{o}(G^{i}(\mathbf{c},\mathbf{a}^{\prime},\bm{\phi}_{i}),\mathbf{s}^{\prime},\bm{\phi}_{o})$
@@
Appendix E Proofs
E.1 Sketch Proof of Theorem 1
Lemma 1.
Given two domains $e_{i},e_{j}\in\mathcal{E}$, $\mathbb{E}_{\mathbb{P}(X^{e_{j}},Z^{e_{j}})}g(f(X^{e_{j}}),Z^{e_{j}})$ can be bounded by $\mathbb{E}_{\mathbb{P}(X^{e_{i}},Z^{e_{i}})}g(f(X^{e_{i}}),Z^{e_{i}})$ as follows:
$\mathbb{E}_{\mathbb{P}(X^{e_{j}},Z^{e_{j}})}g(f(X^{e_{j}}),Z^{e_{j}})\leq\mathbb{E}_{\mathbb{P}(X^{e_{i}},Z^{e_{i}})}g(f(X^{e_{i}}),Z^{e_{i}})+\sqrt{2}D[\mathbb{P}(X^{e_{j}},Z^{e_{j}},Y^{e_{j}}),\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}})]$
Lemma 2.
Given two domains $e_{i},e_{j}\in\mathcal{E}$, under Lemma 1, $\epsilon^{e_{j}}(f)$ can be bounded by $\epsilon^{e_{i}}(f)$ as follows:
$\epsilon^{e_{j}}(f)\leq\epsilon^{e_{i}}(f)+\sqrt{2}D[\mathbb{P}(X^{e_{j}},Z^{e_{j}},Y^{e_{j}}),\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}})]$
Under Lemmas 1 and 2, we now prove Theorem 1
Proof.
Let $e_{\star}\in\mathcal{E}_{s}$ be the source domain nearest to the target domain $e_{t}\in\mathcal{E}\backslash\mathcal{E}_{s}$. Under Lemma 2, we have
$\epsilon^{e_{t}}(f)\leq\epsilon^{e_{i}}(f)+\sqrt{2}D[\mathbb{P}(X^{e_{t}},Z^{e_{t}},Y^{e_{t}}),\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}})]$
where $e_{i}\in\mathcal{E}_{s}$. Taking average of upper bounds based on all source domains, we have:
$\epsilon^{e_{t}}(f)$ $\leq\frac{1}{|\mathcal{E}_{s}|}\sum_{e_{i}\in\mathcal{E}_{s}}\epsilon^{e_{i}}(f)+\frac{\sqrt{2}}{|\mathcal{E}_{s}|}\sum_{e_{i}\in\mathcal{E}_{s}}D[\mathbb{P}(X^{e_{t}},Z^{e_{t}},Y^{e_{t}}),\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}})]$
$\leq\frac{1}{|\mathcal{E}_{s}|}\sum_{e_{i}\in\mathcal{E}_{s}}\epsilon^{e_{i}}(f)+\frac{\sqrt{2}}{|\mathcal{E}_{s}|}|\mathcal{E}_{s}|D[\mathbb{P}(X^{e_{t}},Z^{e_{t}},Y^{e_{t}}),\mathbb{P}(X^{e_{\star}},Z^{e_{\star}},Y^{e_{\star}})]$
$+\frac{\sqrt{2}}{|\mathcal{E}_{s}|}\sum_{e_{i}\in\mathcal{E}_{s}}D[\mathbb{P}(X^{e_{\star}},Z^{e_{\star}},Y^{e_{\star}}),\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}})]$
$\leq\frac{1}{|\mathcal{E}_{s}|}\sum_{e_{i}\in\mathcal{E}_{s}}\epsilon^{e_{i}}(f)+\sqrt{2}\min_{e_{i}\in\mathcal{E}_{s}}D[\mathbb{P}(X^{e_{t}},Z^{e_{t}},Y^{e_{t}}),\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}})]$
$+\sqrt{2}\max_{e_{i},e_{j}\in\mathcal{E}_{s}}D[\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}}),\mathbb{P}(X^{e_{j}},Z^{e_{j}},Y^{e_{j}})]$
E.2 Sketch Proof of Theorem 2
@@
Let $d$ be a distance metric between probability measures for which it holds that $d[\mathbb{P,T}]=0$ for two distributions $\mathbb{P}$ and $\mathbb{T}$ if and only if $\mathbb{P}=\mathbb{T}$ almost surely. Then $P^{\star}(0,0)=P^{\star}$
@@
Assuming the perturbation function $P^{\star}(\gamma_{1},\gamma_{2})$ is $L$-lipschitz continuous in $\gamma_{1},\gamma_{2}$. Then given Proposition 1, it follows that $|P^{\star}-P^{\star}(\gamma_{1},\gamma_{2})|\leq L||\bm{\gamma}||_{1}$, where $\bm{\gamma}=[\gamma_{1},\gamma_{2}]^{T}$.
@@
Let $\Theta\subseteqq\mathbb{R}^{p}$ be a finite-dimensional parameter space. For $\xi>0$, a function $\hat{f}:\mathcal{X\times}\Theta\mathcal{\rightarrow Y}$ is said to be an $\xi$-parameterization of $\mathcal{F}$ if it holds that for each $f\in\mathcal{F}$, there exists a parameter $\bm{\theta}\in\Theta$ such that $\mathbb{E}_{\mathbb{P}(X)}\Arrowvert\hat{f}(\mathbf{x},\bm{\theta})-f(\mathbf{x})\Arrowvert_{\infty}\leq\xi$. Given an $\xi$-parameterization $\hat{f}$ of $\mathcal{F}$, consider the following saddle-point problem:
$D_{\xi}^{\star}(\gamma_{1},\gamma_{2})\triangleq\max_{\lambda_{1}(e_{i},e_{j}),\lambda_{2}(e_{i},e_{j})}\min_{\bm{\theta}\in\Theta}R(\bm{\theta})$ $+\int_{e_{i},e_{j}\in\mathcal{E}}[\delta^{e_{i},e_{j}}(\bm{\theta})-\gamma_{1}]\mathrm{d}\lambda_{1}(e_{i},e_{j})$
$+\int_{e_{i},e_{j}\in\mathcal{E}}[\epsilon^{e_{i}}(\bm{\theta})+\epsilon^{e_{j}}(\bm{\theta})-\gamma_{2}]\mathrm{d}\lambda_{2}(e_{i},e_{j})$
where $R(\bm{\theta})=R(\hat{f}(\cdot,\bm{\theta}))$ and $\mathcal{L}^{e_{i},e_{j}}(\bm{\theta})=\mathcal{L}^{e_{i},e_{j}}(\hat{f}(\cdot,\bm{\theta}))$.
@@
$|d[f_{1}(\mathbf{x}),f_{1}(\mathbf{x}^{\prime}=T(\mathbf{x},z,e))]-d[f_{2}(\mathbf{x}),f_{2}(\mathbf{x}^{\prime}=T(\mathbf{x},z,e))]|\leq L_{d}\Arrowvert f_{1}(\mathbf{x})-f_{2}(\mathbf{x})\Arrowvert_{\infty},\quad\forall e\in\mathcal{E}$
@@
$|(g\circ f_{1})(\mathbf{x},z)-(g\circ f_{2})(\mathbf{x},z)|\leq L_{g}\Arrowvert f_{1}(\mathbf{x})-f_{2}(\mathbf{x})\Arrowvert_{\infty},\quad\forall e\in\mathcal{E}$
@@
There exists a parameter $\bm{\theta}\in\Theta$ such that $\delta^{e_{i},e_{j}}(\bm{\theta})<\gamma_{1}-\xi\cdot\max\{L_{\ell},L_{d}\}$ and $\epsilon^{e_{i}}(\bm{\theta})+\epsilon^{e_{j}}(\bm{\theta})<\gamma_{2}-\xi\cdot\max\{L_{\ell},L_{g}\},\forall e_{i},e_{j}\in\mathcal{E}$
@@
Assume $\ell$ and $d$ are non-negative and bounded in $[-B,B]$ and let $d_{\mathrm{VC}}$ denote the VC-dimension of the hypothesis class $\mathcal{A}_{\xi}=\{\hat{f}(\cdot,\bm{\theta}):\bm{\theta}\in\Theta\}\subseteq\mathcal{F}$. Then it holds with probability $1-\omega$ over the $N$ samples from each domain that
@@
The Theorem 2. Let $\xi>0$ be given, and let $\hat{f}$ be an $\xi$-parameterization of $\mathcal{F}$. Let the assumptions holds, and further assume that $\ell$, $d$, and $g$ are $[0,B]$-bounded and that $d[\mathbb{P,T}]=0$ if and only if $\mathbb{P=T}$ almost surely, and that $P^{\star}(\gamma_{1},\gamma_{2})$ is $L$-Lipschitz. Then assuming that $\mathcal{A}_{\xi}=\{\hat{f}(\cdot,\theta):\theta\in\Theta\}\subseteq\mathcal{F}$ has finite VC-dimension, it holds with probability $1-\omega$ over the $N$ samples that
$|P^{\star}-D^{\star}_{\xi,N,\mathcal{E}_{s}}(\bm{\gamma})|\leq L||\bm{\gamma}||_{1}+\xi k(1+||\bm{\lambda}^{\star}_{p}||_{1})+O(\sqrt{\log(M)/M})$
@@
The proof of this theorem is a simple consequence of the triangle inequality. Indeed, by combining Proposition 2, Proposition 3, and Proposition 4, we find that
@@
Appendix F Additional Results
Additional results including more visualization (Fig. 6), as well as complete results with all domains and baselines on ccMNIST (Tab. 6), FairFace (Tab. 7), FairFace (Tab. 8), and NYSF (Tab. 9) are provided. Additional ablation study results are in Tabs. 10, 11, 12 and 13.
Figure 6: Additional visualizations for data reconstruction and under the transformation model $T$.
@@ Appendix F Additional Results
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$

### v2

@@ Appendix C Ablation Studies
9:         $\mathcal{L}_{inv}(\boldsymbol{\theta})=\mathcal{L}^{\prime}_{inv}(\boldsymbol{\theta})/m$
10:         $\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{cls}(\boldsymbol{\theta})+\lambda_{1}\cdot\mathcal{L}_{inv}(\boldsymbol{\theta})$
11:         $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta_{p}\cdot\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$
12:         $\lambda_{1}\leftarrow\text{max}\{[\lambda_{1}+\eta_{d}\cdot(\mathcal{L}_{inv}(\boldsymbol{\theta})-\gamma_{1})],0\}$
@@
15:procedure T($\mathbf{x},z,y$)
16:     $\mathbf{c}=E^{c}(E^{m}(\mathbf{x},\boldsymbol{\theta}^{m}),\boldsymbol{\theta}^{c})$
@@
19:     $\mathbf{x}^{\prime}=G^{o}(G^{i}(\mathbf{c},\mathbf{a}^{\prime},\boldsymbol{\phi}_{i}),\mathbf{s}^{\prime},\boldsymbol{\phi}_{o})$
@@
Appendix D Proofs
D.1 Sketch Proof of Theorem 1
@@
Let $d^{\prime}$ be a distance metric between probability measures for which it holds that $d^{\prime}[\mathbb{P,T}]=0$ for two distributions $\mathbb{P}$ and $\mathbb{T}$ if and only if $\mathbb{P}=\mathbb{T}$ almost surely. Then $P^{\star}(0,0)=P^{\star}$
@@
Assuming the perturbation function $P^{\star}(\gamma_{1},\gamma_{2})$ is $L$-lipschitz continuous in $\gamma_{1},\gamma_{2}$. Then given Proposition 1, it follows that $|P^{\star}-P^{\star}(\gamma_{1},\gamma_{2})|\leq L||\boldsymbol{\gamma}||_{1}$, where $\boldsymbol{\gamma}=[\gamma_{1},\gamma_{2}]^{T}$.
@@
Let $\Theta\subseteqq\mathbb{R}^{p}$ be a finite-dimensional parameter space. For $\xi>0$, a function $\hat{f}:\mathcal{X}\times\Theta\rightarrow\mathcal{Y}$ is said to be an $\xi$-parameterization of $\mathcal{F}$ if it holds that for each $f\in\mathcal{F}$, there exists a parameter $\boldsymbol{\theta}\in\Theta$ such that $\mathbb{E}_{\mathbb{P}_{X}}\Arrowvert\hat{f}(\mathbf{x},\boldsymbol{\theta})-f(\mathbf{x})\Arrowvert_{\infty}\leq\xi$. Given an $\xi$-parameterization $\hat{f}$ of $\mathcal{F}$, consider the following saddle-point problem:
$D_{\xi}^{\star}(\gamma_{1},\gamma_{2})\triangleq\max_{\lambda_{1}(s_{i},s_{j}),\lambda_{2}(s_{i},s_{j})}\min_{\boldsymbol{\theta}\in\Theta}\>$ $R(\boldsymbol{\theta})+\int_{s_{i},s_{j}\in\mathcal{E}_{s}}[\delta^{s_{i},s_{j}}(\boldsymbol{\theta})-\gamma_{1}]\mathrm{d}\lambda_{1}(s_{i},s_{j})$
$+\int_{s_{i},s_{j}\in\mathcal{E}_{s}}[\rho^{s_{i}}(\boldsymbol{\theta})+\rho^{s_{j}}(\boldsymbol{\theta})-\gamma_{2}]\mathrm{d}\lambda_{2}(s_{i},s_{j})$
where $R(\boldsymbol{\theta})=R(\hat{f}(\cdot,\boldsymbol{\theta}))$ and $\mathcal{L}^{s_{i},s_{j}}(\boldsymbol{\theta})=\mathcal{L}^{s_{i},s_{j}}(\hat{f}(\cdot,\boldsymbol{\theta}))$.
@@
$|d[f_{1}(\mathbf{x}),f_{1}(\mathbf{x}^{\prime}=T(\mathbf{x},z,s))]-d[f_{2}(\mathbf{x}),f_{2}(\mathbf{x}^{\prime}=T(\mathbf{x},z,s))]|\leq L_{d}\Arrowvert f_{1}(\mathbf{x})-f_{2}(\mathbf{x})\Arrowvert_{\infty},\quad\forall s\in\mathcal{E}_{s}$
@@
$|(g\circ f_{1})(\mathbf{x},z)-(g\circ f_{2})(\mathbf{x},z)|\leq L_{g}\Arrowvert f_{1}(\mathbf{x})-f_{2}(\mathbf{x})\Arrowvert_{\infty}$
@@
There exists a parameter $\boldsymbol{\theta}\in\Theta$ such that $\delta^{s_{i},s_{j}}(\boldsymbol{\theta})<\gamma_{1}-\xi\cdot\max\{L_{\ell},L_{d}\}$ and $\rho^{s_{i}}(\boldsymbol{\theta})+\rho^{s_{j}}(\boldsymbol{\theta})<\gamma_{2}-\xi\cdot\max\{L_{\ell},L_{g}\},\forall s_{i},s_{j}\in\mathcal{E}_{s}$
@@
Assume $\ell$ and $d$ are non-negative and bounded in $[-B,B]$ and let $d_{\mathrm{VC}}$ denote the VC-dimension of the hypothesis class $\mathcal{A}_{\xi}=\{\hat{f}(\cdot,\boldsymbol{\theta}):\boldsymbol{\theta}\in\Theta\}\subseteq\mathcal{F}$. Then it holds with probability $1-\omega$ over the $N$ samples from each domain that
@@
Let $\xi>0$ be given, and let $\hat{f}$ be an $\xi$-parameterization of $\mathcal{F}$. Let the assumptions hold, and further assume that $\ell$, $d$, and $g$ are $[0,B]$-bounded and that $d[\mathbb{P,T}]=0$ if and only if $\mathbb{P=T}$ almost surely, and that $P^{\star}(\gamma_{1},\gamma_{2})$ is $L$-Lipschitz. Then assuming that $\mathcal{A}_{\xi}=\{\hat{f}(\cdot,\theta):\theta\in\Theta\}\subseteq\mathcal{F}$ has finite VC-dimension, it holds with probability $1-\omega$ over the $N$ samples that
$|P^{\star}-D^{\star}_{\xi,N,\mathcal{E}_{s}}(\boldsymbol{\gamma})|\leq L||\boldsymbol{\gamma}||_{1}+\xi k(1+||\boldsymbol{\lambda}^{\star}_{p}||_{1})+O(\sqrt{\log(M)/M})$
@@
The proof of this theorem is a simple consequence of the triangle inequality. Indeed, by combining Propositions 2, 3 and 4, we find that
@@
D.2 Sketch Proof of Theorem 2
Lemma 1.
Given two domains $e_{i},e_{j}\in\mathcal{E}$, $\mathbb{E}_{\mathbb{P}^{e_{j}}_{XZ}}g(f(X^{e_{j}}),Z^{e_{j}})$ can be bounded by $\mathbb{E}_{\mathbb{P}^{e_{i}}_{XZ}}g(f(X^{e_{i}}),Z^{e_{i}})$ as follows:
$\mathbb{E}_{\mathbb{P}^{e_{j}}_{XZ}}g(f(X^{e_{j}}),Z^{e_{j}})\leq\mathbb{E}_{\mathbb{P}^{e_{i}}_{XZ}}g(f(X^{e_{i}}),Z^{e_{i}})+\sqrt{2}dist[\mathbb{P}^{e_{j}}_{XZY},\mathbb{P}^{e_{i}}_{XZY}]$
Lemma 2.
Given two domains $e_{i},e_{j}\in\mathcal{E}$, under Lemma 1, $\rho^{e_{j}}(f)$ can be bounded by $\rho^{e_{i}}(f)$ as follows:
$\rho^{e_{j}}(f)\leq\rho^{e_{i}}(f)+\sqrt{2}dist[\mathbb{P}^{e_{j}}_{XZY},\mathbb{P}^{e_{i}}_{XZY}]$
Under Lemmas 1 and 2, we now prove Theorem 2
Proof.
Let $s_{\star}\in\mathcal{E}_{s}$ be the source domain nearest to the target domain $t\in\mathcal{E}\backslash\mathcal{E}_{s}$. Under Lemma 2, we have
$\rho^{t}(f)\leq\rho^{s_{i}}(f)+\sqrt{2}dist[\mathbb{P}^{t}_{XZY},\mathbb{P}^{e_{i}}_{XZY}]$
where $s_{i}\in\mathcal{E}_{s}$. Taking the average of upper bounds based on all source domains, we have:
$\rho^{t}(f)\leq$ $\frac{1}{|\mathcal{E}_{s}|}\sum_{s_{i}\in\mathcal{E}_{s}}\rho^{s_{i}}(f)+\frac{\sqrt{2}}{|\mathcal{E}_{s}|}\sum_{s_{i}\in\mathcal{E}_{s}}dist[\mathbb{P}^{t}_{XZY},\mathbb{P}^{e_{i}}_{XZY}]$
$\leq$ $\frac{1}{|\mathcal{E}_{s}|}\sum_{s_{i}\in\mathcal{E}_{s}}\rho^{s_{i}}(f)+\frac{\sqrt{2}}{|\mathcal{E}_{s}|}|\mathcal{E}_{s}|dist[\mathbb{P}^{t}_{XZY},\mathbb{P}^{s_{\star}}_{XZY}]+\frac{\sqrt{2}}{|\mathcal{E}_{s}|}\sum_{s_{i}\in\mathcal{E}_{s}}dist[\mathbb{P}^{s_{\star}}_{XZY},\mathbb{P}^{s_{i}}_{XZY}]$
$\leq$ $\frac{1}{|\mathcal{E}_{s}|}\sum_{s_{i}\in\mathcal{E}_{s}}\rho^{s_{i}}(f)+\sqrt{2}\min_{s_{i}\in\mathcal{E}_{s}}dist[\mathbb{P}^{t}_{XZY},\mathbb{P}^{e_{i}}_{XZY}]+\sqrt{2}\max_{s_{i},s_{j}\in\mathcal{E}_{s}}dist[\mathbb{P}^{e_{i}}_{XZY},\mathbb{P}^{e_{j}}_{XZY}]$
Appendix E Additional Results
Additional results including complete results with all domains and baselines on ccMNIST (Tab. 7), FairFace (Tab. 8), FairFace (Tab. 9), and NYSF (Tab. 10) are provided. We showcase the reconstruction loss using the FairFace data in Fig. 8. Additional ablation study results are in Tabs. 11, 12, 13 and 14.
@@ Appendix E Additional Results
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$